# thanks: corresponding author

### v2

@@ Abstract
We construct a model for the nucleation of a wormhole within a Lorentzian spacetime by employing techniques from topological surgery and Morse theory. In our framework, a 0-surgery process describes the neighborhood of the nucleation point inside a compact region of spacetime, yielding a singular Lorentzian cobordism that connects two spacelike regions with different topologies. To avoid the singularity at the critical point of the Morse function, we employ the Misner trick of taking a connected sum with a closed $4$-manifold—namely $\mathbb{CP}^{2}$—to obtain an everywhere non‑degenerate Lorentzian metric. This connected sum replaces the naked singularity with a region containing closed timelike curves. The obtained spacetime is non-singular, but violates all the standard energy conditions. Our construction thus shows that a wormhole can be “created” without singularities in classical general relativity.
@@ I.1 Motivation
Two fundamental perspectives emerge in addressing classical topology changes. The first prioritizes causality and leads to scenarios where the interpolating spacetimes include regions with metric degeneracies [8, 9], Morse singularities [10, 11, 12], or instantons [13, 14, 15, 16, 17, 18, 19]. The second emphasizes the equivalence principle and permits topology changes at the expense of causality [20, 21], leading to conceptual issues such as causality paradoxes [22, 23, 24].
@@ II Topological surgery
In Figure 1, the level sets $f<0$ are diffeomorphic to each other, as are the level sets $f>0$. However, between them there is the critical level set: the double cone $f=0$.
@@ III.1 Preliminaries
As reported by Geroch [7], the first proof that two closed $3$-manifolds are always cobordant was due to Rohklin [56], using cobordism theory. In fact, two closed manifolds are cobordant if and only if their Stiefel-Whitney numbers are the same [48]. If, in addition, the manifolds are oriented, then their Pontryagin numbers must also be equal. However, all closed $3$-manifolds have the same Pontryagin and Euler numbers. For more details on the subject, please see Refs. [48, 57, 58].
@@
Thus a Lorentzian cobordism between two closed spacelike $3$-manifolds always exists, provided the Euler characteristic is adjusted by an appropriate connected sum. This cobordism, as imposed by the second Geroch theorem [7], must however contain either singularities or CTCs if $\Sigma_{i}$ and $\Sigma_{f}$ are not diffeomorphic:
@@
Let $W$ be a compact spacetime whose boundary is the disjoint union of two compact spacelike $3$-manifolds, $\Sigma_{i}$ and $\Sigma_{f}$. If $W$ has no CTCs, then $\Sigma_{i}$ and $\Sigma_{f}$ are diffeomorphic, and further $W$ is topologically equivalent to $\Sigma_{i}\times[0,1]$.
@@
As discussed in Ref. [60], the presence of CTCs and the consequent violations of causality may be physically more troubling than the mere lack of predictability stemming from singularities. Indeed, a theorem due to Hawking demonstrates that the formation of singularities is driven by energy conditions rather than by chronological assumptions, which are usually invoked in other singularity theorems [^2] See Theorem 4, p. 272 of Ref. [60]. This result is strengthened by a series of theorems due to Tipler [63, 64] and Lee [65].
In particular, the formation of singularities in topology-changing spacetimes—even in our compact-transition setup—appears unavoidable if the energy conditions hold [^3] See Theorem 4, p. 20 of Ref. [63]. Moreover, even the mere existence of a compact, non-simply-connected region in spacetime forces violations of the energy conditions, as demonstrated by a theorem due to Gannon [66]. Because classical matter generally satisfies the energy conditions, it therefore seems unlikely that a wormhole can nucleate under an evolution which is nearly classical [67].
@@ III.2 The kink number
$\operatorname{kink}(\Sigma;V^{\mu})\equiv\sum_{i}^{m}\mathrm{sign}(\chi_{i}),$
where $\chi_{i}$ are the intersection points and $\mathrm{sign}(\chi_{i})=+1$ iff the orientation of $S(\Sigma)$ at $\chi_{i}$ equals the product of the orientations of $V^{\mu}$ and $N^{\mu}$, and 1-1 otherwise.
Because $\Sigma$ is orientable and $3$-dimensional, one can choose a global frame $\{(e^{\mu}_{0}=N^{\mu},e^{\mu}_{i})\leavevmode\nobreak\ |\leavevmode\nobreak\ i=1,2,3\}$ on $\Sigma$, where $e^{\mu}_{0}$ is the inward unit normal and $e_{i}^{\mu}$ are tangent. Extending this frame to a collared neighborhood $C\cong\Sigma\times\left[0,1\right]$, the decomposition
@@
$\operatorname{kink}(\Sigma;V^{\mu})\equiv\mathrm{deg}(f^{a}).$
@@
$J\equiv\mathrm{det}\begin{pmatrix}f^{0}&f^{1}&f^{2}&f^{3}\\ D_{1}f^{0}&D_{1}f^{1}&D_{1}f^{2}&D_{1}f^{3}\\ D_{2}f^{0}&D_{2}f^{1}&D_{2}f^{2}&D_{2}f^{3}\\ D_{3}f^{0}&D_{3}f^{1}&D_{3}f^{2}&D_{3}f^{3}\end{pmatrix}$, (5)
@@
$\operatorname{kink}(\partial W;V^{\mu})=\sum_{j=1}^{m}\leavevmode\nobreak\ \operatorname{kink}(\Sigma_{j};V^{\mu}).$
@@ III.3 Trivial 2+1 cobordisms
$\int_{\theta_{-}}^{\theta_{+}}J\,\mathrm{d}\theta=\int_{\theta_{-}}^{\theta_{+}}\frac{g(\theta)\leavevmode\nobreak\ g^{\prime}(\theta)}{\left(1-2g(\theta)+2g^{2}(\theta)\right)^{\frac{3}{2}}}\,\mathrm{d}\theta=1.$ (9)
@@ III.4 The wormhole nucleation cobordism
$x_{1}^{2}+x_{2}^{2}+\cdots+x_{n}^{2}=r_{0}^{2},$
in an $n$-dimensional Minkowski spacetime is equal to one. Moreover, as anticipated, the contribution to the kink number only comes from the regions of $\partial W$ where the vector field changes from timelike to spacelike and vice versa; it is therefore independent of the interior topology of $\Sigma_{i}$ and $\Sigma_{f}$.
@@ IV.1 Introduction
Taking the connected sum $W=\mathbb{M}\#\mathbb{CP}^{2}$ requires the removal of a 4-ball $D_{1}^{4}$ around the critical point $p\in\mathbb{M}$ of the Morse function, leaving a boundary $3$-sphere $V_{1}$. Similarly, we remove a $4$-ball $D_{2}^{4}$ from $\mathbb{CP}^{2}$ around the unique isolated singular point $q$ of a chosen vector field $w^{\mu}$, yielding another boundary $3$-sphere $V_{2}$. The identification of these boundaries is achieved via an orientation-reversing gluing diffeomorphism $h:V_{1}\to V_{2}$, ensuring that the resulting space $W$ inherits a topological-manifold structure through the quotient topology [47].
@@ IV.2 The Morse spacetime
$g_{L}^{\mu\nu}=g_{R}^{\mu\nu}+2\frac{\partial^{\mu}f\partial^{\nu}f}{g_{R}^{\alpha\beta}\partial_{\alpha}f\partial_{\beta}f},$
@@
$\mathrm{d}s^{2}_{\mathrm{M}}=\frac{-U\leavevmode\nobreak\ (z\,\mathrm{d}z-r\,\mathrm{d}r)^{2}+(r\,\mathrm{d}z+z\,\mathrm{d}r)^{2}}{z^{2}+r^{2}}+r^{2}\mathrm{d}\Omega_{2}^{2},$ (13)
@@
$\mathrm{d}s^{2}_{\text{E}}=\frac{\mathrm{d}\rho^{2}}{2\sqrt{t^{2}+\rho^{2}}}+R(t,\rho)\,\mathrm{d}\phi^{2}.$
@@
$\mathrm{d}s^{2}=\mathrm{d}z^{2}+\mathrm{d}r^{2}+r^{2}\leavevmode\nobreak\ \mathrm{d}\phi^{2},$
@@
$\left[1+\left(\frac{\mathrm{d}z}{\mathrm{d}r}\right)^{2}\right]=\frac{2(r^{2}-t_{c})}{r^{2}-2t_{c}}.$
@@
For timelike geodesics we set $\epsilon=-1$ and solve numerically. Choosing the initial conditions $\rho(0)=5$ and $t(0)=1$, the initial velocity must satisfy
$(1+2\sqrt{26})\leavevmode\nobreak\ \dot{t}^{2}_{0}+\dot{\rho}^{2}_{0}=-2\sqrt{26}.$
@@
Figure 7: Numerical solutions of Eqs. (15a)–(15b) for timelike and lightlike geodesics passing through the wormhole. The integration is performed over the affine interval $\lambda\in[0,10]$ with $t(0)=1$ and $\rho(0)=5$.
@@
where $l^{\mu}_{+}$ and $l^{\mu}_{-}$ are real null vectors and $m^{\mu},\bar{m}^{\mu}$ are complex-conjugate null vectors. Using this null tetrad, one can deduce that the Morse spacetime is of Petrov type D.
@@
$\mathrm{d}s^{2}_{\mathrm{M}}=(\gamma_{\mu\nu}-l^{+}_{\mu}l^{-}_{\nu}-l^{-}_{\mu}l^{+}_{\nu})\leavevmode\nobreak\ \mathrm{d}x^{\mu}\mathrm{d}x^{\nu}.$
@@
$A\left(\Sigma_{\pm}\right)=\int_{\Sigma_{\pm}}\sqrt{\gamma}\leavevmode\nobreak\ \mathrm{d}^{2}x,$
@@
$\left.\frac{\mathrm{d}\theta_{\pm}}{\mathrm{d}u_{\pm}}\right|_{\theta_{\pm}=0}=\frac{At^{2}\pm t\rho\sqrt{A}+\rho^{2}\sqrt{t^{2}+\rho^{2}}}{A^{2}R\leavevmode\nobreak\ (t^{2}+\rho^{2})},$
@@
$\mathbf{e}_{z}$ $=-\sqrt{\frac{C}{F}}\leavevmode\nobreak\ \mathrm{d}z+\frac{zr(F+2)}{\sqrt{CF}}\leavevmode\nobreak\ \mathrm{d}r,$ (18a)
$\mathbf{e}_{r}$ $=\sqrt{\frac{F(F+1)}{C}}\leavevmode\nobreak\ \mathrm{d}r,$ (18b)
$\mathbf{e}_{\theta}$ $=r\leavevmode\nobreak\ \mathrm{d}\theta,\qquad\mathbf{e}_{\phi}=r\sin(\theta)\leavevmode\nobreak\ \mathrm{d}\phi.$ (18c)
@@
Finally, the regions highlighted in Fig. 9 are also those where the Segre type changes from $[ZZ(11)]$ (type I) to $[(1,1)11]$ (type IV); on the boundary the Segre type is $[(2,11)]$.
@@ IV.3 A singular Lorentzian metric on the complex projective plane
enables us to deduce, by Proposition 7.1 of Ref. [87], that the line at infinity $\mathbb{CP}^{1}_{\infty}$ is not a leaf of the singular holomorphic foliation.
@@
The absence of symmetries makes the analysis of this spacetime challenging; accordingly, we focus solely on the behavior of timelike geodesic congruences. Specifically, we choose a particular family of geodesic congruences, which we integrate numerically. The initial conditions for the congruence are eleven random points on the unit $3$-sphere centered at the origin of the $U_{1}$ chart, so that
@@
Figure 11: Numerical solution of the geodesic equations for the metric (22) obtained via a Fehlberg fourth-fifth-order Runge-Kutta scheme. Initial conditions were chosen as eleven random points on the unit 33-sphere, with the initial velocity given by the (normalized) vector field (21). (a): Projection of the trajectories onto the xyzxyz hyperplane. (b): Projection of the trajectories onto the xytxyt hyperplane.
@@ V.2 Interface between the complex projective plane and the annulus
$\mathrm{d}s^{2}_{\text{FS}}=\mathrm{d}\chi^{2}+A^{2}(\chi)\left(\bm{\sigma}_{1}^{2}+\bm{\sigma}_{2}^{2}+B^{2}(\chi)\leavevmode\nobreak\ \bm{\sigma}_{3}^{2}\right),$ (30)
@@
$\bm{\sigma}_{1}$ $=\cos\psi\leavevmode\nobreak\ \mathrm{d}\theta+\sin\psi\sin\theta\leavevmode\nobreak\ \mathrm{d}\phi,$
$\bm{\sigma}_{2}$ $=-\sin\psi\leavevmode\nobreak\ \mathrm{d}\theta+\cos\psi\sin\theta\leavevmode\nobreak\ \mathrm{d}\phi,$
$\bm{\sigma}_{3}$ $=\mathrm{d}\psi+\cos\theta\leavevmode\nobreak\ \mathrm{d}\phi,$
@@
$\mathrm{d}s^{2}_{\mathrm{N}}=\mathrm{d}\tau^{2}+a(\tau)^{2}\,\left(\bm{\sigma}_{1}^{2}+\bm{\sigma}_{2}^{2}+(1-b^{2}(\tau))\leavevmode\nobreak\ \bm{\sigma}_{3}^{2}\right).$ (31)
@@
$\left[\frac{\partial}{\partial x^{\alpha}}g_{\mu\nu,\alpha}\right]=0,$ (32)

### v3

@@ Abstract
We construct a model for the nucleation of a wormhole within a Lorentzian spacetime by employing techniques from topological surgery and Morse theory. In our framework, a 0-surgery process describes the neighborhood of the nucleation point inside a compact region of spacetime, yielding a singular Lorentzian cobordism that connects two spacelike regions with different topologies. To avoid the singularity at the critical point of the Morse function, we employ the Misner trick of taking a connected sum with a closed 4-manifold—namely $\mathbb{CP}^{2}$—to obtain an everywhere nondegenerate Lorentzian metric. This connected sum replaces the naked singularity with a region containing closed timelike curves. The obtained spacetime is nonsingular, but violates all the standard energy conditions. Our construction, thus, shows that a wormhole can be “created” without singularities in classical general relativity.
@@ I.1 Motivation
Two fundamental perspectives emerge in addressing classical topology changes. The first prioritizes causality and leads to scenarios where the interpolating spacetimes include regions with metric degeneracies [8, 9], Morse singularities [10, 11, 12], or instantons [13, 14, 15, 16, 17, 18]. The second emphasizes the equivalence principle and permits topology changes at the expense of causality [19, 20], leading to conceptual issues such as causality paradoxes [21, 22, 23].
@@ II Topological surgery
Here and in the rest of this paper, $S^{n}$ represent topological spheres and $D^{n}$ represent closed disks.
@@
In Fig. 1, the level sets $f<0$ are diffeomorphic to each other, as are the level sets $f>0$. However, between them there is the critical level set: the double cone $f=0$.
@@ III.1 Preliminaries
As reported by Geroch, two closed 3-manifolds are always cobordant [7]. In fact, two closed manifolds are cobordant if and only if their Stiefel-Whitney numbers are the same [47]. If, in addition, the manifolds are oriented, then their Pontryagin numbers must also be equal. However, all closed $3$-manifolds have the same Pontryagin and Euler numbers (in particular, Pontryagin numbers for odd-dimensional manifolds are always zero). For more details on the subject, please see Refs. [47, 55, 56].
@@
Thus a Lorentzian cobordism between two closed spacelike $3$-manifolds always exists, provided the Euler characteristic is adjusted by an appropriate connected sum. This cobordism, as imposed by the second Geroch theorem [7], must, however contain either singularities or CTCs if $\Sigma_{i}$ and $\Sigma_{f}$ are not diffeomorphic:
@@
Let $W$ be a compact spacetime whose boundary is the disjoint union of two compact spacelike $3$-manifolds, $\Sigma_{i}$ and $\Sigma_{f}$. If $W$ has no CTCs, then $\Sigma_{i}$ and $\Sigma_{f}$ are diffeomorphic, and, furthermore, $W$ is topologically equivalent to $\Sigma_{i}\times[0,1]$.
@@
As discussed in Ref. [58], the presence of CTCs and the consequent violations of causality may be physically more troubling than the mere lack of predictability stemming from singularities. Indeed, a theorem due to Hawking demonstrates that the formation of singularities is driven by energy conditions rather than by chronological assumptions, which are usually invoked in other singularity theorems [^2] See Theorem 4, p. 272 in Ref. [58]. This result is strengthened by a series of theorems due to Tipler [61, 62] and Lee [63].
In particular, the formation of singularities in topology-changing spacetimes—even in our compact-transition setup—appears unavoidable if the energy conditions hold [^3] See Theorem 4, p. 20 in Ref. [61]. Moreover, even the mere existence of a compact, non-simply-connected region in spacetime forces violations of the energy conditions, as demonstrated by a theorem due to Gannon [64]. Because classical matter generally satisfies the energy conditions, it therefore seems unlikely that a wormhole can nucleate under an evolution which is nearly classical [65].
@@ III.2 The kink number
$\operatorname{kink}(\Sigma;V^{\mu})\equiv\sum_{i}^{m}\mathrm{sgn}(\chi_{i}),$
where $\chi_{i}$ are the intersection points and $\mathrm{sgn}(\chi_{i})=+1$ iff the orientation of $S(\Sigma)$ at $\chi_{i}$ equals the product of the orientations of $V^{\mu}$ and $N^{\mu}$, and 1-1 otherwise.
Because $\Sigma$ is orientable and $3$-dimensional, one can choose a global frame $\{(e^{\mu}_{0}=N^{\mu},e^{\mu}_{i})\penalty 10000\ |\penalty 10000\ i=1,2,3\}$ on $\Sigma$, where $e^{\mu}_{0}$ is the inward unit normal and $e_{i}^{\mu}$ are tangent. Extending this frame to a collared neighborhood $C\cong\Sigma\times\left[0,1\right]$, the decomposition
@@
$\operatorname{kink}(\Sigma;V^{\mu})\equiv\deg(f^{a}).$
@@
$J\equiv\det\begin{pmatrix}f^{0}&f^{1}&f^{2}&f^{3}\\ D_{1}f^{0}&D_{1}f^{1}&D_{1}f^{2}&D_{1}f^{3}\\ D_{2}f^{0}&D_{2}f^{1}&D_{2}f^{2}&D_{2}f^{3}\\ D_{3}f^{0}&D_{3}f^{1}&D_{3}f^{2}&D_{3}f^{3}\end{pmatrix}$, (5)
@@
$\operatorname{kink}(\partial W;V^{\mu})=\sum_{j=1}^{m}\penalty 10000\ \operatorname{kink}(\Sigma_{j};V^{\mu}).$
@@ III.3 Trivial 2+1 cobordisms
$\int_{\theta_{-}}^{\theta_{+}}J\,\mathrm{d}\theta=\int_{\theta_{-}}^{\theta_{+}}\frac{g(\theta)\penalty 10000\ g^{\prime}(\theta)}{\left(1-2g(\theta)+2g^{2}(\theta)\right)^{\frac{3}{2}}}\,\mathrm{d}\theta=1.$ (9)
@@ III.4 The wormhole nucleation cobordism
$x_{1}^{2}+x_{2}^{2}+\cdots+x_{n}^{2}=r_{0}^{2}$
in an $n$-dimensional Minkowski spacetime is equal to one. Moreover, as anticipated, the contribution to the kink number comes only from the regions of $\partial W$ where the vector field changes from timelike to spacelike and vice versa; it is therefore independent of the interior topology of $\Sigma_{i}$ and $\Sigma_{f}$.
@@ IV.1 Introduction
Taking the connected sum $W=\mathbb{M}\#\mathbb{CP}^{2}$ requires the removal of a 4-ball $B_{1}^{4}$ around the critical point $p\in\mathbb{M}$ of the Morse function, leaving a boundary $3$-sphere $V_{1}$. Similarly, we remove a $4$-ball $B_{2}^{4}$ from $\mathbb{CP}^{2}$ around the unique isolated singular point $q$ of a chosen vector field $w^{\mu}$, yielding another boundary $3$-sphere $V_{2}$. The identification of these boundaries is achieved via an orientation-reversing gluing diffeomorphism $h:V_{1}\to V_{2}$, ensuring that the resulting space $W$ inherits a topological-manifold structure through the quotient topology [46].
@@ IV.2 The Morse spacetime
$g_{L}^{\mu\nu}=g_{R}^{\mu\alpha}g_{R}^{\nu\beta}g^{L}_{\alpha\beta},$
@@
$\mathrm{d}s^{2}_{\mathrm{M}}=\frac{-U\penalty 10000\ (z\,\mathrm{d}z-r\,\mathrm{d}r)^{2}+(r\,\mathrm{d}z+z\,\mathrm{d}r)^{2}}{z^{2}+r^{2}}+r^{2}\mathrm{d}\Omega_{2}^{2},$ (13)
@@
$\mathrm{d}s^{2}_{\text{E}}=\frac{\mathrm{d}\rho^{2}}{2\sqrt{t_{c}^{2}+\rho^{2}}}+R(t_{c},\rho)\,\mathrm{d}\phi^{2}.$
@@
$\mathrm{d}s^{2}=\mathrm{d}z^{2}+\mathrm{d}r^{2}+r^{2}\penalty 10000\ \mathrm{d}\phi^{2},$
@@
$\left[1+\left(\frac{\mathrm{d}z}{\mathrm{d}r}\right)^{2}\right]=\frac{2(r^{2}-t_{c})}{r^{2}-2t_{c}},$
@@
For timelike geodesics, we set $\epsilon=-1$ and solve numerically. Choosing the initial conditions $\rho(0)=5$ and $t(0)=1$, the initial velocity must satisfy
$(1+2\sqrt{26})\penalty 10000\ \dot{t}^{2}_{0}+\dot{\rho}^{2}_{0}=-2\sqrt{26}.$
@@
Figure 7: Numerical solutions of Eqs. (15a) and (15b) for timelike and lightlike geodesics passing through the wormhole. The integration is performed over the affine interval $\lambda\in[0,10]$ with $t(0)=1$ and $\rho(0)=5$.
@@
where $l^{\mu}_{+}$ and $l^{\mu}_{-}$ are real null vectors and $m^{\mu}$ and $\bar{m}^{\mu}$ are complex-conjugate null vectors. Using this null tetrad, one can deduce that the Morse spacetime is of Petrov type D.
@@
$\mathrm{d}s^{2}_{\mathrm{M}}=(\gamma_{\mu\nu}-l^{+}_{\mu}l^{-}_{\nu}-l^{-}_{\mu}l^{+}_{\nu})\penalty 10000\ \mathrm{d}x^{\mu}\mathrm{d}x^{\nu}.$
@@
$A\left(\Sigma_{\pm}\right)=\int_{\Sigma_{\pm}}\sqrt{\gamma}\penalty 10000\ \mathrm{d}^{2}x,$
@@
$\left.\frac{\mathrm{d}\theta_{\pm}}{\mathrm{d}u_{\pm}}\right|_{\theta_{\pm}=0}=\frac{At^{2}\pm t\rho\sqrt{A}+\rho^{2}\sqrt{t^{2}+\rho^{2}}}{A^{2}R\penalty 10000\ (t^{2}+\rho^{2})},$
@@
$\mathbf{e}_{z}$ $=-\sqrt{\frac{C}{F}}\,\mathrm{d}z+\frac{zr(F+2)}{\sqrt{CF}}\,\mathrm{d}r,$ (18a)
$\mathbf{e}_{r}$ $=\sqrt{\frac{F(F+1)}{C}}\,\mathrm{d}r,$ (18b)
$\mathbf{e}_{\theta}$ $=r\,\mathrm{d}\theta,\qquad\mathbf{e}_{\phi}=r\sin(\theta)\,\mathrm{d}\phi.$ (18c)
@@
Finally, the regions highlighted in Fig. 9 are also those where the Segre type changes from $Z\bar{Z}(11)]$ (type IV) to $[(1,1)11]$ (type I); on the boundary the Segre type is $[2,(11)]$.
@@ IV.3 A singular Lorentzian metric on the complex projective plane
enables us to deduce, by Proposition 7.1 in Ref. [85], that the line at infinity $\mathbb{CP}^{1}_{\infty}$ is not a leaf of the singular holomorphic foliation.
@@
The absence of symmetries makes the analysis of this spacetime challenging; accordingly, we focus solely on the behavior of timelike geodesic congruences. Specifically, we choose a particular family of geodesic congruences, which we integrate numerically. The initial conditions for the congruence are 11 random points on the unit $3$-sphere centered at the origin of the $U_{1}$ chart, so that
@@
Figure 11: Numerical solution of the geodesic equations for the metric (22) obtained via a Fehlberg fourth-fifth-order Runge-Kutta scheme. Initial conditions were chosen as 11 random points on the unit 33-sphere, with the initial velocity given by the (normalized) vector field (21). (a): Projection of the trajectories onto the xyzxyz hyperplane. (b): Projection of the trajectories onto the xytxyt hyperplane.
@@ V.2 Interface between the complex projective plane and the annulus
$\mathrm{d}s^{2}_{\text{FS}}=\mathrm{d}\chi^{2}+A^{2}(\chi)\left(\bm{\sigma}_{1}^{2}+\bm{\sigma}_{2}^{2}+B^{2}(\chi)\penalty 10000\ \bm{\sigma}_{3}^{2}\right),$ (30)
@@
$\bm{\sigma}_{1}$ $=\cos\psi\penalty 10000\ \mathrm{d}\theta+\sin\psi\sin\theta\penalty 10000\ \mathrm{d}\phi,$
$\bm{\sigma}_{2}$ $=-\sin\psi\penalty 10000\ \mathrm{d}\theta+\cos\psi\sin\theta\penalty 10000\ \mathrm{d}\phi,$
$\bm{\sigma}_{3}$ $=\mathrm{d}\psi+\cos\theta\penalty 10000\ \mathrm{d}\phi,$
@@
$\mathrm{d}s^{2}_{\mathrm{N}}=\mathrm{d}\tau^{2}+a(\tau)^{2}\,\left(\bm{\sigma}_{1}^{2}+\bm{\sigma}_{2}^{2}+(1-b^{2}(\tau))\penalty 10000\ \bm{\sigma}_{3}^{2}\right).$ (31)
@@
$\left[g_{\mu\nu,\alpha}\right]=0,$ (32)